OpenStudy (anonymous):

Please help!!! (algebra) If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. What is the numerator of the fraction?

OpenStudy (anonymous):

thanks for stopping by! can you help me?

OpenStudy (anonymous):

Let n=numerator and d=denominator a) (n+3)/d=3/4 b) n/(d+7)=1 b) n=d+7 a) (d+7+3)/d=3/4 substitute d+7 for n. a) (d+10)/d=3/4

OpenStudy (anonymous):

a) 4(d+10)=3d multiply each side by 4d a) 4d+40=3d d=-40 b) n=-40+7 n=-33 .

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Okay thank you, seiga!

OpenStudy (anonymous):

np.

mathslover (mathslover):

@seiga Check the question again, the denominator is decreased by 7 NOT increased So we have : \(\large{\frac{n}{d-7} = 1}\) n = d-7 d = n + 7 ----------(1)

OpenStudy (anonymous):

i see...

mathslover (mathslover):

Now... since \(\large{\frac{n+3}{d} = \frac{3}{4}}\) \(\large{4(n+3) = 3d}\) \(\large{4n+12 = 3d}\) Put d = n+7 4n + 12 = 3(n+7) 4n+ 12 = 3n +21 n = 9 So we have numerator = 9 and d = n + 7 = 16 hence numerator = 9 and denominator = 16

mathslover (mathslover):

Got it now @scienceQs ?

OpenStudy (anonymous):

yes! thank you =)